OpenStudy (briellshaw):

Find the table that follows this rule: multiply by 5 and subtract 2. Input 3 5 7 10 12 Output 13 23 33 48 58 Input 3 5 7 10 12 Output 15 25 35 50 60 Input 3 5 7 10 12 Output 13 23 33 52 62 Input 3 5 7 10 12 Output 13 23 33 44 54

OpenStudy (alphadxg):

@BriellShaw Well, we need to look at the input, take the input of any number and use the rule "5 x input number" = A Then take "A - 2" = output number

OpenStudy (alphadxg):

@BriellShaw Every single number in the input coloumn has to follow the "5 X input number rule" then you will get "A". then Take "A - 2" this will give you your output number

OpenStudy (anonymous):

exactly as @alphadxg says are there any options u can eliminate ?

OpenStudy (briellshaw):

The answer is B

OpenStudy (briellshaw):

thats the wron answer

OpenStudy (alphadxg):

The answer should be A, because (input) 3) 3x5 = 15 -2 = 13 (output) (input) 5) 5x5 = 25 -2 = 23 (output) (input) 7) 7x5 = 35 -2 = 33 (output) (input) 10) 10 x 5 = 50 - 2 = 48 (output) (input) 12) 12x5 = 60 - 2 = 58 (output)